If we have a initial concentration of H4NO2 (aq) at .10 M, make an ICE table for the following reaction, and solve for the concentration of NH4, NO2, NH3 and HNO2

NH4+(aq)+NO2-(aq)<=>NH3(aq)+HNO2(aq)

(not sure if this helps, but the value of K1 is 1.23e-6)
Answer #1

-4 . [NH₂] = [HNO2) = x= 1.89 x 10 9 (M) [NHst] = [NO] = 0.1-4 = 0.09989 (M)